OpenStudy (anonymous):

x-int = -1/9 y-int = 9 put in point slope form and slope intercept form

OpenStudy (anonymous):

I got y-9 for the first side

OpenStudy (anonymous):

but i dont know where to go from there

OpenStudy (anonymous):

Y=-X/9+9 this is the answer

OpenStudy (anonymous):

Webwork says its wrong

OpenStudy (anonymous):

then maybe Y=-1/9*X+9

OpenStudy (anonymous):

nope, says it cant be negative. tried positive and it says its wrong to

OpenStudy (anonymous):

here, ill take a screenshot

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i cant look at it idk why.

OpenStudy (anonymous):

hmm well it says to put it in point slope form first so y-y1=m(x-x1) so i got y-9 for the first side which was right, but when i put m(x-(-1/9)) it says it was wrong

OpenStudy (anonymous):

i dont know how to find the slope

OpenStudy (anonymous):

try -9. otherwise idk what to do.

OpenStudy (anonymous):

nope, idk either